# Neural network learns low-dimensional polynomials with SGD near the information-theoretic limit

### v1

@@ B.3.1 When $\mathrm{IE}(\sigma_{*}^{I})=2$ with $I\geq 2$
$=(\eta d)^{k}\bigg{|}\sum_{l=0}^{k+1}{k+1\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{k}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{k-l+1}\mathsf{He}_{j-k-1}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.48)
$=(\eta d)^{k}\bigg{|}\sum_{l=0}^{k+1}{k+1\choose l}\mathbb{E}[\upsilon^{l}]\sum_{m=3}^{\infty}m!\kappa^{q(k-l+1)}H\big{(}(\sigma_{*})^{k-l+1};m\big{)}H\big{(}(\sigma^{\prime})^{k}\mathsf{He}_{j-k-1};m-1\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}\bigg{|}$ (B.49)
$\lesssim(\eta d)^{k}\kappa^{2}.$ (B.50)
When $k\geq I$, we know that $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{k+1},\cdots,(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I}$ have information exponent larger than $1$, and $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I-1},\cdots,(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))$ have information exponent larger than $2$. Thus, the expectation in (B.46) is bounded by
[…]
$=(\eta d)^{k}\bigg{|}\sum_{l=0}^{k+1}{k+1\choose l}\mathbb{E}[\upsilon^{l}]\sum_{m=2}^{\infty}m!\kappa^{q(k-l+1)}H\big{(}(\sigma_{*})^{k-l+1};m-1\big{)}H\big{(}(\sigma^{\prime})^{k}\mathsf{He}_{j-k-1};m\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}\bigg{|}$ (B.52)
$\lesssim(\eta d)^{k}\kappa.$ (B.53)
Now, consider the case when $k=I-1$:
$\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}\mathbb{E}\bigg{[}y\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})+\upsilon\big{)}\big{)}^{I}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.54)
$=\sum_{l=0}^{I}\sum_{j=1}^{C_{\sigma}}(\eta d)^{k}j\beta_{j}{j-1\choose I-1}{I\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}.$ (B.55)
Note that $\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}),\cdots,(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I-1}$ has information exponent larger than $2$. Therefore, for $l\geq 1$, we have
$\bigg{|}\boldsymbol{\theta}^{\top}(\eta d)^{I-1}j\beta_{j}{j-1\choose I-1}{I\choose l}\mathbb{E}[\upsilon^{l}]\mathbb{E}\bigg{[}\mathsf{He}_{j-1-k}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.56)
$\lesssim(\eta d)^{I-1}\bigg{|}\mathbb{E}\bigg{[}\sum_{m=3}^{\infty}m!\kappa^{m-1}H\big{(}\mathsf{He}_{j-I}(\sigma^{\prime})^{I-1};m-1\big{)}H\big{(}\big{(}\sigma_{*}\big{)}^{I-l};m\big{)}\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}\bigg{|}$ (B.57)
$\lesssim(\eta d)^{I-1}\kappa^{2}.$ (B.58)
And for $l=0$, $(\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x}))^{I}$ has information exponent $2$ and we have
$\sum_{j=1}^{C_{\sigma}}(\eta d)^{I-1}j\beta_{j}{j-1\choose I-1}\mathbb{E}\bigg{[}\mathsf{He}_{j-I}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}\big{)}\big{(}\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})\big{)}^{I-1}\big{(}\sigma_{*}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\big{)}^{I-l}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}\bigg{]}$ (B.59)
$=(\eta d)^{I-1}\sum_{m=2}^{\infty}m!\kappa^{m-1}H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};m-1\bigg{)}H\big{(}(\sigma_{*})^{I};m\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}$ (B.60)
If $H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};1\bigg{)}\neq 0$, we have
(B.60)$=(1+O(\kappa))\cdot(\eta d)^{I-1}\underbrace{2H\bigg{(}\sum_{j=1}^{C_{\sigma}}j\beta_{j}{j-1\choose I-1}(\sigma^{\prime})^{I-1}\mathsf{He}_{j-I};1\bigg{)}H\big{(}(\sigma_{*})^{I};2\big{)}}_{=:\gamma}\kappa\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{v}.$ (B.61)
We have that
$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]$ (B.62)
$=(1+O(\kappa))\cdot(\eta d)^{I-1}\gamma\kappa\boldsymbol{\theta}^{\top}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}+O(\kappa^{2}+C_{1}\eta d^{\frac{1}{2}}+(\eta d)^{I-2}\kappa^{2}+(\eta d)^{I}\kappa+(\eta d)^{I-1}\kappa^{2})$ (B.63)
When $d^{-\frac{1}{2}}C_{3}\leq\kappa\leq c_{2}$ and $\eta=c_{1}d^{-1}$, we have
$\boldsymbol{\theta}^{\top}\mathbb{E}[y\sigma^{\prime}\big{(}\boldsymbol{\omega}^{\top}\boldsymbol{x}+\eta d\sigma^{\prime}(\boldsymbol{\omega}^{\top}\boldsymbol{x})y\big{)}\boldsymbol{P}_{\boldsymbol{\omega}}\boldsymbol{x}]=(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa$ (B.64)
Together with the bound on the first term ((B.37) and (B.38)), the expectation of $\boldsymbol{g}$ (B.25) is evaluated as
$\boldsymbol{\theta}^{\top}\mathbb{E}[\boldsymbol{g}]=(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa.$ (B.65)
By using this, the expected update of the alignment becomes
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ $\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\mathbb{E}[\|\boldsymbol{g}\|^{2}]-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}].$ (B.66)
Note that $\mathbb{E}[\|\boldsymbol{g}\|^{2}],\mathbb{E}[|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}]\lesssim d$ and $\kappa\lesssim c_{1}$. Thus,
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa-C_{1}\eta\tilde{\eta}^{2}(\kappa\eta d+\eta^{2}\tilde{\eta}d).$ (B.67)
When $\tilde{\eta}\leq c_{2}$, $\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]$ is evaluated as
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]\geq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa.$ (B.68)
In the same way, using (B.29), we also have the opposite bound:
$\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]\leq\kappa+\eta\tilde{\eta}(1+O(c_{1}))\cdot c_{1}^{I-1}\gamma\kappa$ (B.69)
Regarding the noise, recall that
$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\frac{\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}}{\|\boldsymbol{\omega}+\eta\tilde{\eta}\boldsymbol{g}\|}.$ (B.70)
$\eta^{-1}\tilde{\eta}^{-1}\big{(}\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\mathbb{E}[\kappa+\eta\tilde{\eta}\boldsymbol{\theta}^{\top}\boldsymbol{g}]\big{)}=\ \boldsymbol{\theta}^{\top}\boldsymbol{g}-\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{g}]$ is a mean-zero polynomial of Gaussian inputs, with all coefficients and variances of inputs bounded by $O(1)$. Notice that normalization does not increase the absolute value of the noise. Thus, regarding $\nu=\eta^{-1}\tilde{\eta}^{-1}\big{(}\boldsymbol{\theta}^{\top}\boldsymbol{w}-\mathbb{E}[\boldsymbol{\theta}^{\top}\boldsymbol{w}]\big{)}$, we have that
$\mathbb{P}\big{[}\big{|}\eta^{-1}\tilde{\eta}^{-1}\boldsymbol{\theta}^{\top}\boldsymbol{w}-\mathbb{E}[\eta^{-1}\tilde{\eta}^{-1}\boldsymbol{\theta}^{\top}\boldsymbol{w}]\big{|}>t\big{]}\leq\mathbb{P}\bigg{[}\bigg{|}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\mathbb{E}\bigg{[}\boldsymbol{\theta}^{\top}\boldsymbol{g}\bigg{]}\bigg{|}>t\bigg{]}\leq\exp(-t^{1/C_{1}}/C_{1}).$ (B.71)
This completes the proof. ∎
B.3.2 When $\mathrm{IE}(\sigma_{*}^{I})=1$ with $I\geq 2$
[…]
@@ B.3.2 When $\mathrm{IE}(\sigma_{*}^{I})=1$ with $I\geq 2$
For the case of $\mathrm{IE}(\sigma_{*}^{I})=1$, starting from $\boldsymbol{w}=\boldsymbol{\omega}$, if we choose step size $\eta=c_{a}\eta^{t}=c_{1}d^{-1}$ and negative momentum $\xi=1-\tilde{\eta}$, and assume that $d^{-\frac{1}{2}}\leq\kappa=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}\leq c_{2}$ and $\tilde{\eta}\leq c_{2}$, then the expected change in the alignment after two gradient steps on the same sample $(\boldsymbol{x},y)$ in Algorithm 1 is as follows:
$\boldsymbol{\theta}^{\top}\boldsymbol{w}=\boldsymbol{\theta}^{\top}\boldsymbol{\omega}+(1+O(c_{1}))\cdot\eta\tilde{\eta}c_{1}^{I-1}H\bigg{(}\frac{1}{(I-1)!}\sigma^{(I)}(\sigma^{(1)})^{I-1};0\bigg{)}H\big{(}(\sigma_{*})^{I};1\big{)}+\eta\tilde{\eta}\nu,$ (B.72)
where $\nu$ is a mean-zero random variable that satisfies $\mathbb{P}[|\nu|>s]\leq\exp(-s^{1/C_{1}}/C_{1})$ for all $s>0$.
Proof.  Similarly to Lemma 9, update of the alignment is evaluated as
$\boldsymbol{\theta}^{\top}\boldsymbol{w}\geq\kappa+\eta c_{2}\boldsymbol{\theta}^{\top}\boldsymbol{g}-\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2}$ (B.73)
and
$\boldsymbol{\theta}^{\top}\boldsymbol{w}\leq\kappa+\eta c_{2}\boldsymbol{\theta}^{\top}\boldsymbol{g}+\frac{1}{2}\kappa\eta^{2}\tilde{\eta}^{2}\|\boldsymbol{g}\|^{2}+\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2},$ (B.74)

### v2

@@ B.4 Weak Recovery: Stochastic Update
when $\tau\leq c_{1}^{-1}c_{\xi}^{-1}d-1$. By letting $c_{\xi}\leq c_{\eta}^{-4}c_{1}C_{1}^{-1}\delta$, we have $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq c_{\eta}c_{\xi}^{\frac{1}{2}}c_{1}^{-\frac{1}{2}}C_{1}^{\frac{1}{2}}\delta^{-\frac{1}{2}}d^{-\frac{1}{2}}$, and
$\kappa^{2(\tau+1)}\geq c_{\eta}^{-1}d^{-\frac{1}{2}}+c_{1}c_{\xi}(\tau+1)d^{-1},$ (B.82)
which verifies $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq\kappa^{2t}$ for $t=\tau+1$. Thus, there exists some $\tau_{*}\leq c_{1}^{-1}c_{\xi}^{-1}d$ such that
$\kappa^{2\tau_{*}}\geq 4c_{2},$ (B.83)
for $c_{1}\leq\frac{1}{4}c_{\eta}^{I}$, with probability $1-\delta$.
Now we prove that $\kappa^{2t}\geq 2c_{2}$ holds for all $\tau_{*}\leq t\leq T_{1,1}=C_{3}c_{\xi}^{-1}d$. Because $\nu^{2t}$ are mean-zero sub-Weibull random variables, we also have that $|\sum_{s=\tau}^{\tau+\tau^{\prime}-1}\nu^{2s}|\leq C_{4}\sqrt{\tau^{\prime}}$ for all $0\leq\tau,\tau^{\prime}\leq T_{1,1}$ with high probability. Also, because $\eta^{t}\ll d^{-1}$ and $|1-\xi^{t}|\ll 1$, we can easily see that $|\kappa^{2(\tau+1)}-\kappa^{2\tau}|=\tilde{O}(d^{-1})$ for all $\tau=0,1,\cdots,T_{1,1}-1$, with high probability. Thus, when there exists $\tau\geq\tau_{*}$ such that $\kappa^{2(\tau-1)}\geq 4c_{2}$ and $\kappa^{2\tau}<4c_{2}$, we have $\kappa^{2\tau}\geq 3c_{2}$ with high probability. Moreover, following the above argument, we can inductively show that
$\kappa^{2(\tau+\tau^{\prime})}$ $\geq 3c_{2}+c_{1}c_{\xi}\tau^{\prime}d^{-1}-c_{\eta}c_{\xi}d^{-1}C_{4}\sqrt{\tau^{\prime}}$ (B.84)
$\geq 3c_{2}+c_{1}c_{\xi}\tau^{\prime}d^{-1}-\begin{cases}c_{2}&(\tau^{\prime}\leq c_{\eta}^{-2}c_{\xi}^{-2}C_{4}^{-2}c_{2}^{2}d^{2})\\ c_{1}c_{\xi}\tau^{\prime}d^{-1}&(\tau^{\prime}\geq c_{\eta}^{2}c_{1}^{-2}C_{4}^{2})\end{cases}.$ (B.85)
$\geq 2c_{2},$ (B.86)
for $\tau^{\prime}\leq T_{1,1}=C_{3}c_{\xi}^{-1}d$ or until $\kappa^{2(\tau+\tau^{\prime})}\geq 4c_{2}$ holds again. By repeating this argument (if there are multiple such $\tau$), we see that $\kappa^{2t}\geq 2c_{2}$ holds for all $\tau_{*}\leq t\leq T_{1,1}=C_{3}c_{\xi}^{-1}d$ with high probability.
(ii) When $p_{*}=2$. We define $\iota_{0}=0,\iota_{1}=\log_{(1+c_{1}c_{\xi}d^{-1})}(4),\iota_{2}=2\log_{(1+c_{1}c_{\xi}d^{-1})}(4),\dots$. We show that, for each $i$, if $\kappa^{2\iota_{i}}\geq 2c_{\eta}^{-1}d^{-\frac{1}{2}}$, we have $\kappa^{2(\iota_{i+1})}\geq 2\kappa^{2\iota_{i}}$, with probability at least $1-\delta 4^{-i}$, or there exists some $t\ (\iota_{i}<t\leq\iota_{i+1})$ with $\kappa^{2t}>c_{\eta}^{I}$.
Assume that the above statement holds until some $i-1\geq 0$ (we do not need to assume anything for $i=0$). Then, we have $\kappa^{2\iota_{i}}\geq 2^{i}\kappa^{0}\geq 2c_{\eta}^{-1}d^{-\frac{1}{2}}$. Similarly to (B.80), if $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq\kappa^{2t}\leq c_{\eta}^{I}$ for all $t=\iota_{i},\iota_{i}+1,\cdots,\tau$, we have
$\kappa^{2(\tau+1)}\geq\kappa^{2\iota_{i}}+c_{1}c_{\xi}d^{-1}\sum_{s=\iota_{i}}^{\tau}\kappa^{2s}-c_{\eta}c_{\xi}d^{-1}\bigg{|}\sum_{s=\iota_{i}}^{\tau}\nu^{2s}\bigg{|}.$ (B.87)
Applying (B.78) with $\delta=\delta/4^{i}$ and $T=\frac{1}{4}c_{\eta}^{-2}c_{\xi}^{-2}C^{-1}(\delta/4^{i})(\kappa^{2\iota_{i}})^{2}d^{2}-1$ to get
$\kappa^{2(\tau+1)}$ $\geq\kappa^{2\iota_{i}}+c_{1}c_{\xi}d^{-1}\sum_{s=\iota_{i}}^{\tau}\kappa^{2s}-c_{\eta}c_{\xi}d^{-1}C^{\frac{1}{2}}\delta^{-\frac{1}{2}}\sqrt{\tau+1-\iota_{i}}$ (B.88)
$\geq\kappa^{2\iota_{i}}+c_{1}c_{\xi}d^{-1}\sum_{s=\iota_{i}}^{\tau}\kappa^{2s}-\frac{1}{2}\kappa^{2\iota_{i}}$ (B.89)
when $\tau\leq\iota_{i}+\frac{1}{4}c_{\eta}^{-2}c_{\xi}^{-2}C^{-1}(\delta/4^{i})(\kappa^{2\iota_{i}})^{2}d^{2}-1$, which verifies $c_{\eta}^{-1}d^{-\frac{1}{2}}\leq\frac{1}{2}\kappa^{2\iota_{i}}\leq\kappa^{2t}$ for $t=\tau+1$.
This implies that, with probability $1-\delta/4^{i}$, we have
$\kappa^{2(\tau+1)}\geq\frac{1}{2}\kappa^{2\iota_{i}}+c_{1}c_{\xi}d^{-1}\sum_{s=\iota_{i}}^{\tau}\kappa^{2s}$ (B.90)
for all $\tau=\frac{1}{4}c_{\eta}^{-2}c_{\xi}^{-2}C^{-1}(\delta/4^{i})(\kappa^{2\iota_{i}})^{2}d^{2}-1$, which is equivalent to
$\kappa^{2\tau}\geq(1+c_{1}c_{\xi}d^{-1})^{\tau-\iota_{i}}\frac{1}{2}\kappa^{2\iota_{i}}$ (B.91)
for all $\tau=\iota_{i},\iota_{i}+1,\cdots,\iota_{i}+\frac{1}{4}c_{\eta}^{-2}c_{\xi}^{-2}C^{-1}(\delta/4^{i})(\kappa^{2\iota_{i}})^{2}d^{2}$. By taking $c_{\xi}\ll c_{1}c_{\eta}^{-2}C^{-1}(\delta/4^{i})(\kappa^{2\iota_{i}})^{2}d$, we have $\frac{1}{4}c_{\eta}^{-2}c_{\xi}^{-2}C^{-1}(\delta/4^{i})(\kappa^{2\iota_{i}})^{2}d^{2}\geq\log_{(1+c_{1}c_{\xi}d^{-1})}(4)$, and we get
$\kappa^{2\iota_{i+1}}\geq 2\kappa^{2\iota_{i}}$ (B.92)
with probability $1-\delta/4^{i}$ (or there exists $t\leq\iota_{i+1}$ such that $\kappa^{2t}>c_{\eta}^{I}$).
Thus, by induction, for all $i$, we have that
$\kappa^{2\iota_{i}}\geq 2^{i}\kappa^{0},$ (B.93)
or that there exists some $t\leq\iota_{i}$ such that $\kappa^{2t}$ is larger than $c_{\eta}^{I}$, with probability $1-\delta$.
The LHS of (B.93) becomes larger than $c_{\eta}^{I}$ for some $i\leq\log d$. Because $\iota_{i}=\Theta(ic_{1}^{-1}c_{\xi}^{-1}d)$, within $O(c_{1}^{-1}c_{\xi}d\log d)$ steps, there exists at least one $\tau_{*}=O(c_{1}^{-1}c_{\xi}^{-1}d\log d)$ such that $\kappa^{2\tau_{*}}\geq 4c_{2}$ for $c_{2}\leq\frac{1}{4}c_{\eta}^{I}$, with probability $1-\delta$.
Once such $\tau_{*}$ is obtained, following the last paragraph of (i), we can see that $\kappa^{2t}\geq 2c_{2}$ holds until $t=T_{1,1}$ with high probability.
(iii) When $p_{*}\geq 3$. We apply (B.78) with $T=\frac{1}{p_{*}-2}c_{1}^{-1}c_{\xi}^{-1}d^{\frac{p_{*}}{2}}(\kappa^{0})^{-(p_{*}-2)}$ to obtain that
$c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}}\bigg{|}\sum_{s=0}^{\tau}\nu^{2s}\bigg{|}\leq c_{\eta}c_{\xi}^{\frac{1}{2}}c_{1}^{-\frac{1}{2}}C^{\frac{1}{2}}\delta^{-\frac{1}{2}}d^{-\frac{p_{*}}{4}}(\kappa^{0})^{-\frac{p_{*}-2}{2}}$ (B.94)
for all $\tau=0,1,\cdots,T-1$, with probability $1-\delta$.
We take $c_{\xi}\ll c_{\eta}^{-2}c_{1}C^{-1}\delta d^{\frac{p_{*}}{4}}(\kappa^{0})^{\frac{p_{*}}{2}}$ so that (B.94) is bounded by $c_{1}^{-1}d^{-\frac{1}{2}}$. Then,
$\kappa^{2(\tau+1)}$ $\geq\kappa^{0}+c_{1}c_{\xi}d^{-\frac{p_{*}}{2}}\sum_{s=0}^{\tau}(\kappa^{2s})^{p_{*}-1}+c_{\eta}c_{\xi}d^{-\frac{p_{*}}{2}}\sum_{s=0}^{\tau}\nu^{2s}$ (B.95)
$\geq c_{\eta}^{-1}d^{-\frac{1}{2}}+c_{1}c_{\xi}d^{-\frac{p_{*}}{2}}\sum_{s=0}^{\tau}(\kappa^{2\tau})^{p_{*}-1}.$ (B.96)
It is easy to see that $\kappa^{2(\tau+1)}$ is lower bounded by $a^{\tau+1}$, where $a^{0}=c_{\eta}^{-1}d^{-\frac{1}{2}}$ and $a^{\tau+1}=a^{\tau}+c_{1}c_{\xi}d^{-\frac{p_{*}}{2}}(a^{\tau})^{p_{*}-1}$. By applying Lemma 18, we have
$\kappa^{2\tau}\geq\frac{\kappa^{0}}{\mathopen{}\mathclose{{}\left(1-c_{1}c_{\xi}d^{-\frac{p_{*}}{2}}(p_{*}-2)(\kappa^{0})^{(p_{*}-2)}t}\right)^{\frac{1}{p_{*}-2}}}.$ (B.97)
Thus, until $\tau\leq\big{(}c_{1}c_{\xi}d^{-\frac{p_{*}}{2}}(p_{*}-2)(\kappa^{0})^{(p_{*}-2)}\big{)}^{-1}\leq T+1\ll d^{p_{*}-1}$, with probability at least $1-\delta$, there exists some $\tau_{*}$ such that
$\kappa^{2\tau_{*}}\geq 4c_{2}\geq c_{\eta}^{I}$ (B.98)
when $c_{2}\leq\frac{1}{4}c_{\eta}^{I}$.
Once such $\tau_{*}$ is obtained, following the last paragraph of (i), we can see that $\kappa^{2t}\geq 2c_{2}$ holds until $t=T_{1,1}$ with high probability. ∎
In the above proof we used the (discrete version of) Bihari–LaSalle inequality from [BAGJ22].
@@
For $p\geq 3$ and $c>0$, consider a positive sequence $(a^{t})_{t\geq 0}$ such that
$a^{t+1}=a^{t}+c(a^{t})^{p-1}.$ (B.99)
Then, we have
$a^{t}\geq\frac{a^{0}}{\mathopen{}\mathclose{{}\left(1-c(p-2)(a^{0})^{(p-2)}t}\right)^{\frac{1}{p-2}}}.$ (B.100)
Proof.  From definition, we have
$c=\frac{a^{t+1}-a^{t}}{(a^{t})^{p-1}}\leq\int_{t=a^{t}}^{a^{t+1}}\frac{1}{x^{p-1}}\leq\frac{1}{p-2}\mathopen{}\mathclose{{}\left[\frac{1}{(a^{t})^{p-2}}-\frac{1}{(a^{t+1})^{p-2}}}\right].$ (B.101)
Taking the summation and re-arranging the terms yield
$(a^{t})^{-(p-2)}\leq(a^{0})^{-(p-2)}-c(p-2)t,$ (B.102)
$\therefore a^{t}\geq\frac{a^{0}}{\mathopen{}\mathclose{{}\left(1-c(p-2)(a^{0})^{(p-2)}t}\right)^{\frac{1}{p-2}}},$ (B.103)
which gives the lower bound. ∎
B.5 From Weak Recovery to Strong Recovery
In the previous subsection, we proved that after $t=2T_{1,1}=\tilde{\Theta}(d)$ steps, with probability $\tilde{\Omega}(1)$ over the randomness of initialization, we obtain nontrivial alignment $\kappa^{2T_{1,1}}_{j}\geq 2c_{2}$. This subsection discusses how to convert the weak recovery into the strong recovery.
Lemma 19.
Suppose the neuron satisfies $\kappa^{2T_{1,1}}\geq 2c_{2}$. Take $\eta^{2t}=\eta=\bar{c}_{\eta}\varepsilon d^{-1}$, $\eta^{2t+1}=0$, $\xi^{2(t+1)}=0$ for all $t\geq T_{1,1}$, where $\bar{c}_{\eta}\lesssim\mathrm{poly}(c_{1})$. If the activation functions satisfy all of the assumptions in Section B.1 for strong recovery, then we have
$\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(T_{1,1}+\tau_{*})}\geq 1-\varepsilon,$ (B.104)
with high probability, where $\tau_{*}\leq T_{1,2}=C_{3}d\varepsilon^{-2}$. Moreover, $\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2(T_{1,1}+t)}\geq 1-\varepsilon$ for all $\tau_{*}\leq t\leq T_{1,2}=C_{3}d\varepsilon^{-2}$, with high probability.
Proof.  Consider the Hermite expansions of $\sigma_{*}$ and $\sigma$. Let $p$ be the smallest degree that both $\sigma_{*}$ and $\sigma$ have non-zero coefficients. First we compute the population gradient (of the correlation term) as
$\mathbb{E}\big{[}\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\big{]}$ $=\mathbb{E}\bigg{[}\tilde{\nabla}_{\boldsymbol{w}}\bigg{(}\sum_{j=p}^{\infty}\alpha_{j}\mathsf{He}_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{x})\bigg{)}\bigg{(}\sum_{j=0}^{\infty}\beta_{j}\mathsf{He}_{j}({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\bigg{)}\bigg{]}$ (B.105)
$=\sum_{j=p}^{\infty}\big{[}j!\alpha_{j}\beta_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j-1}\boldsymbol{\theta}+(j+2)!\alpha_{j}\beta_{j+2}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j}\boldsymbol{w}^{2t}\big{]}.$ (B.106)
Applying $P_{\boldsymbol{w}^{2t}}$, we have
$\mathbb{E}\big{[}P_{\boldsymbol{w}^{2t}}\tilde{\nabla}_{\boldsymbol{w}}y\sigma({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{x})\big{]}=(\boldsymbol{\theta}-({\boldsymbol{w}^{2t}}^{\top}\boldsymbol{\theta})\boldsymbol{w}^{2t})\sum_{j=p}^{\infty}j!\alpha_{j}\beta_{j}(\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t})^{j-1}.$ (B.107)
Thus, the update of the alignment $\kappa^{2t}=\boldsymbol{\theta}^{\top}\boldsymbol{w}^{2t}$ is
$\kappa^{2(t+1)}\geq\kappa^{2t}+\eta\boldsymbol{\theta}^{\top}\boldsymbol{g}-\frac{1}{2}\eta^{2}\kappa^{2t}\|\boldsymbol{g}\|^{2}-\frac{1}{2}\eta^{3}\tilde{\eta}^{3}|\boldsymbol{\theta}^{\top}\boldsymbol{g}|\|\boldsymbol{g}\|^{2},$ (B.108)